OpenStudy (trixitang):

Find the diameter of a cone that has a volume of 104.67 cubic inches and a height of 4 inches. Use 3.14 for pi. (1 point) 4 inches 6 inches 8 inches 10 inches

pooja195 (pooja195):

\[v=πr^{2}\frac{ h }{ 3 }\]

pooja195 (pooja195):

We are given \[104.67=(3.14)r^{2}\frac{ 4 }{ 3 }\] solve for r

OpenStudy (trixitang):

okay so i got 3.14 x 4/3

pooja195 (pooja195):

Good what does that equal?

OpenStudy (trixitang):

1.333333

OpenStudy (trixitang):

sooo 4.18=1.33333 r^2

pooja195 (pooja195):

it should have been 104.67=4.186667r^2

OpenStudy (trixitang):

ohhh what did i do wrong ?

pooja195 (pooja195):

Multiplied wrong.

OpenStudy (trixitang):

ohh i see

pooja195 (pooja195):

Continue to solve for r

pooja195 (pooja195):

Need help?

OpenStudy (trixitang):

yes

pooja195 (pooja195):

Rmemeber we do all the subtracting Subtract 4.186667r^2 from both sides. 104.67−4.186667^r2=4.186667r^2−4.186667r2 −4.186667r2+104.67=0 Subtract 104.67 from both sides. −4.186667r^2+104.67−104.67=0−104.67 −4.186667r^2=−104.67 Now we have \[\huge~\rm~\frac{ −4.186667r^2=−104.67 }{ −4.186667 }\]

OpenStudy (trixitang):

so I just divide the both ?

pooja195 (pooja195):

Divide both sides

OpenStudy (trixitang):

okay give me a sec

OpenStudy (trixitang):

okay so i cancelled the other one out but was left with -108.856667 for the other

pooja195 (pooja195):

That's incorrect. a negative divided by a negative results in a postitive what is -104.67/−4.186667=?

OpenStudy (trixitang):

so 108.856667 r^2 ?

pooja195 (pooja195):

I used a calculator and got r^2=25.000794

OpenStudy (trixitang):

maybe im dividing incorrectly, okay so what do I do now ?

pooja195 (pooja195):

Square root both sides \[\huge~\rm~\sqrt{r^2 }=\sqrt{25}\]

pooja195 (pooja195):

r=?

OpenStudy (trixitang):

i dont know how to find the radius because i confused myself earlier with diameter and stuff

pooja195 (pooja195):

Its basically algerbra. If you know the formula all you do is plug in the given info and isolate the vairble. In this case we need to square root both sides to isolate the "r^2" So what is the square root of 25?

OpenStudy (trixitang):

5

pooja195 (pooja195):

Good so r=5 Radius times 2=diameter so 5 x 2=?

OpenStudy (trixitang):

10

OpenStudy (trixitang):

thank you ! So can you check my answer for this ? Find the height of a square pyramid that has a volume of 32 cubic feet and a base length of 4 feet. (1 point) 2 feet 4 feet 6 feet 8 feet --------------------------------------------------------- i put 6 feet, am i correct ?

pooja195 (pooja195):

Yep that's perfect well done

OpenStudy (trixitang):

yay thank you ! If i have any more questions, ill tag you !

pooja195 (pooja195):

haha you're welcome. And alright feel free to do so! :D

OpenStudy (katie_rogers101):

I'm glad you learned how to do this! If your'e ever in a quick fix with questions like this google calculator offers solutions to geometry problems. Type in solve for a cone, enter in the data requested and poof theres your answer.